# Unveiling the Unexplored Decay Mode of a Light Charged Higgs Boson to an Off-Shell Top Quark and a Bottom Quark

### v1

@@ V Conclusions
In this study, we explored an unexplored decay channel of the light charged Higgs boson into an off-shell top quark and a bottom quark, $H^{\pm}\rightarrow t^{*}b$, within the context of the type-I two-Higgs-doublet model. We focused on the mass range of 130 to 170 GeV, where the decay $H^{\pm}\rightarrow t^{*}b$ becomes significant. To probe this new physics signal without resorting to specific model parameters, we proposed the pair production of charged Higgs bosons as a golden production channel, followed by the decay $H^{+}H^{-}\rightarrow t^{*}b\tau\nu\rightarrow bbjj\tau\nu$.
@@
We conducted a detailed signal-to-background analysis at the High-Luminosity Large Hadron Collider and a prospective 100 TeV proton-proton collider, employing comprehensive cut-flow strategies and the Boosted Decision Tree method. However, we found that the signal significance remains far below the threshold for a confident detection at these colliders, primarily due to the inherent softness of the $b$ jets in the decay process, which fail to meet the basic threshold for jet clustering.
@@
Recognizing the constraints of hadron colliders, we extended our analysis to explore the discovery potential of a multi-TeV muon collider (MuC) for the $H^{+}H^{-}\rightarrow t^{*}b\tau\nu\rightarrow bbjj\tau\nu$ signal process. The MuC offers a significant advantage by fully exploiting the beam energy in collisions between fundamental particles. Our analysis demonstrated that the MuC, particularly at a center-of-mass energy of 3 TeV, provides a promising environment for probing the $H^{\pm}\rightarrow t^{*}b$ decay mode. The cut-flow analysis at the 3 TeV MuC yielded a high signal significance, surpassing the $5\sigma$ discovery threshold, with a total integrated luminosity of 1 ab-1. Specifically, for $M_{H^{\pm}}=130$, 150, and 170 GeV, the signal significances were 13.7, 13.5, and 6.06, respectively. In contrast, the 10 TeV MuC, despite its higher collision energy, requires a substantially larger integrated luminosity of 10 ab-1 to achieve comparable results. This is due to the reduced signal cross sections of the Drell-Yan production process at higher collision energies.
@@
The results of our study underscore the challenges and potential in searching for the light charged Higgs boson via the $t^{*}b$ decay mode. While high-energy hadron colliders face significant obstacles due to the soft $b$ jets, the multi-TeV MuC emerges as a highly effective platform. The detailed simulation and cut-flow strategy developed in this work provide a robust framework for future experimental searches for this promising new signal, emphasizing the critical role of a multi-TeV MuC in exploring beyond the Standard Model physics.
Acknowledgements.

### v2

@@ 5 Conclusions
In this study, we presented the first comprehensive analysis investigating the discovery potential of future colliders for the decay channel of the light charged Higgs boson into an off-shell top quark and a bottom quark, $H^{\pm}\rightarrow t^{*}b$, within the context of the Type-I two-Higgs-doublet model. We focused on the mass range of 130 to 170 GeV, where the $H^{\pm}\rightarrow t^{*}b$ decay becomes particularly significant. To probe this new physics signal without introducing dependence on specific model parameters, we identified pair production of charged Higgs bosons as a robust production channel, followed by the decay $H^{+}H^{-}\rightarrow t^{*}b\tau\nu\rightarrow bbjj\tau\nu$.
@@
We conducted a detailed signal-to-background analysis at the HL-LHC and a prospective 100 TeV proton-proton collider, employing both comprehensive cut-flow strategies and the Boosted Decision Tree (BDT) method at the detector level. However, the results showed that the signal significance remains well below the threshold for confident detection at these colliders, primarily due to the inherent softness of the $b$ jets in the decay process, which struggle to meet basic jet clustering thresholds.
@@
Given these constraints at hadron colliders, we extended our analysis to explore the discovery potential of a multi-TeV muon collider (MuC) for the $H^{+}H^{-}\rightarrow t^{*}b\tau\nu\rightarrow bbjj\tau\nu$ signal process. The MuC offers significant advantages by fully exploiting the beam energy in fundamental particle collisions. Our analysis demonstrated that a MuC, particularly at a center-of-mass energy of 3 TeV, provides a promising environment for probing the $H^{\pm}\rightarrow t^{*}b$ decay mode. The cut-flow analysis at the 3 TeV MuC achieved a high signal significance, surpassing the $5\sigma$ discovery threshold with a total integrated luminosity of 1 ab-1. Specifically, for $M_{H^{\pm}}=130$, 150, and 170 GeV, the signal significances are 13.7, 13.5, and 6.06, respectively. In contrast, the 10 TeV MuC, despite its higher collision energy, requires a substantially larger integrated luminosity of 10 ab-1 to achieve comparable results due to the reduced signal cross sections of the Drell-Yan production process at higher energies.
@@
The results of our study underscore both the challenges and opportunities in searching for the light charged Higgs boson via the $t^{*}b$ decay mode. While high-energy hadron colliders face significant obstacles due to the soft $b$ jets, the multi-TeV MuC emerges as a highly effective platform. The detailed simulation and cut-flow strategy developed in this work provide a robust framework for future experimental searches for this promising new signal, emphasizing the critical role of a multi-TeV MuC in advancing beyond the Standard Model physics.
Acknowledgements
@@
Appendix A Feynman Diagram Analysis of ${\mu^{+}\mu^{-}}\to H^{+}H^{-}\nu\bar{\nu}$
Figure 9: Representative Feynman diagrams for the process ${\mu^{+}\mu^{-}}\to H^{+}H^{-}\nu\bar{\nu}$, which significantly contribute to the total cross section.
In this appendix, we discuss the process:
$\mu^{+}\mu^{-}\to H^{+}H^{-}\nu\bar{\nu},$ (45)
where $\nu$ includes all three neutrino flavors. As a $2\to 4$ scattering process, this reaction involves 61 distinct Feynman diagrams in the Higgs alignment limit of the Type-I 2HDM. In counting these diagrams, we have neglected the Yukawa couplings involving $H^{\pm}$-$e^{\pm}$-$\nu_{e}$ and $H^{\pm}$-$\mu^{\pm}$-$\nu_{\mu}$, given their negligible contributions. This large number of diagrams complicates the analysis of how the total cross section depends on the model parameters.
To efficiently compute the contributions of individual Feynman diagrams to the total cross section, as well as their interference effects, we used CalcHEP Belyaev:2012qa , which is well-suited for this type of analysis. We found that six classes of diagrams provide significant contributions, while the remaining diagrams have negligible effects. The dominant Feynman diagrams are illustrated in Figure 9: (a) Vector boson scattering (VBS) via $W^{+}W^{-}\rightarrow h^{*}\rightarrow H^{+}H^{-}$; (b) VBS via $t$-channel exchange of $H$ and $A$; (c) VBS via $W^{+}W^{-}\rightarrow Z^{/}\gamma^{\rightarrow}H^{+}H^{-}$; (d) Direct VBS contribution from $W^{+}W^{-}\rightarrow H^{+}H^{-}$; (e) $Z^{/}\gamma^{\to}H^{+}H^{-}$ radiation from the initial $\mu^{-}$ or $\mu^{+}$; (f) $Z^{*}\to H^{+}H^{-}$ radiation from the final-state $\nu_{\mu}$ or $\bar{\nu}_{\mu}$.
Among these, diagram (b) exhibits significant sensitivity to $M_{H/A}$. However, this diagram strongly interferes destructively with diagrams (c) and (d) across the allowed parameter space, resulting in only a weak dependence of the total cross section on $M_{H/A}$. Additionally, diagrams (e) and (f) also interfere destructively with diagrams (b), (c), and (d). For our benchmark scenario with $M_{H}=M_{A}=200{\;{\rm GeV}}$, $\tan\beta=10$, and $m_{12}^{2}=3.84\times 10^{3}{\;{\rm GeV}}^{2}$, diagram (a) dominates due to minimal interference effects. Nevertheless, for other parameter choices, interference significantly alters the relative contributions of these diagrams, making it challenging to identify a universally dominant Feynman diagram.